# MemEvo: Memory-Evolving Incremental Multi-view Clustering

### v1

@@ Introduction
Model optimization
We solve Eq.(10) using the Alternating Direction Method of Multipliers (ADMM) (Boyd et al. 2011). It decomposes the original problem into easily solvable subproblems by introducing auxiliary variables and Lagrange multipliers. To facilitate the solution, we introduce auxiliary variables $\mathbf{E}_{t}$ and $\mathcal{M}$ instead of $\mathbf{X}_{t}-\mathbf{Z}_{t}\mathbf{A}_{t}$ and $\mathcal{Z}$, respectively. Then, the Lagrangian function can be written as
@@
$+\frac{\mu}{2}\|\mathbf{X}_{t}-\mathbf{Z}_{t}\mathbf{A}_{t}-\mathbf{E}_{t}+\frac{\mathbf{Y}_{t}}{\mu}\|_{F}^{2}+\frac{\rho}{2}\|\mathcal{Z}-\mathcal{M}+\frac{\mathcal{J}}{\rho}\|_{F}^{2}.$
where $\mathbf{Y}$ and $\mathcal{J}$ denotes the Lagrange multiplier, $\mu>0,\rho>0$ is the penalty parameter.
$\triangleright$ Initial view processing ($t=1$).
For the first view, no history information exists, so the objective function simplifies to the base view representation learning part:
@@
Update $\mathbf{A}_{1}$.
Fixing $\mathbf{Z}_{1}$, $\mathbf{E}_{1}$, we can obtain $\mathbf{A}_{1}$ by:
@@
This is a standard Orthogonal Procrustes Problem (Wen et al. 2018). Let $\mathbf{B}=\mu\mathbf{X}_{1}-\mu\mathbf{E}_{1}+\mathbf{Y}_{1}$, by performing a singular value decomposition (SVD) of $\mathbf{B}$ to obtain $\mathbf{B}=\mathbf{U}_{A}\mathbf{S}\mathbf{V}_{A}^{T}$, then the closed solution of $\mathbf{A}_{1}$ is $\mathbf{A}_{1}=\mathbf{U}_{A}\mathbf{V}_{A}^{T}$. $\mathbf{U}_{A}$ and $\mathbf{V}_{A}$ are the left and right singular value matrices of $\mathbf{B}\mathbf{Z}_{1}^{T}$.
Update $\mathbf{Z}_{1}$.
Fixing the other variables, the subproblem of $\mathbf{Z}_{1}$ can be solved in the following way
@@
Update $\mathbf{E}_{1}$.
Fixing other variables, $\mathbf{E}_{1}$ can be obtained by
@@
The solution of Eq.(LABEL:E) can be obtained by minimizing the thresholding operator in (Zhang et al. 2015), which is
${(\mathbf{E}_{1})}_{(:,i)}=\left\{\begin{matrix}\frac{\|{(\mathbf{C}_{1})}_{(:,i)}\|_{2}-\frac{1}{\mu}}{\|{(\mathbf{C}_{1})}_{(:,i)}\|_{2}}{(\mathbf{C}_{1})}_{(:,i)}&if\ \|{(\mathbf{C}_{1})}_{(:,i)}\|_{2}>\frac{1}{\mu};\\ 0&Otherwise.\end{matrix}\right$. (17)
where $\mathbf{C}_{1}=\mathbf{X}_{1}-\mathbf{Z}_{1}\mathbf{A}_{1}+\frac{\mathbf{Y}_{1}}{\mu}$
Update $\mathbf{Y}_{1}$ and $\mu$.
The Lagrange multipliers and penalty parameters can be updated by the following rules
$\mathbf{Y}_{1}=\mathbf{Y}_{1}+\mu(\mathbf{X}_{1}-\mathbf{Z}_{1}\mathbf{A}_{1}-\mathbf{E}_{1}),\mu=min(\delta\mu,\mu_{max})$ (18)
where $\delta>1$ to accelerate convergence, $\mu_{max}$ are the positive constants. $\mathbf{Z}_{1}$ will serve as the starting point and first historical memory for the subsequent optimization process.
$\triangleright$ Updates to subsequent views ($t>1$).
Update $\mathbf{A}_{t}$ and $\mathbf{E}_{t}$.
Similarly, the update rules for $\mathbf{A}_{t}$ and $\mathbf{E}_{t}$ are similar to the first view from Eq.(13) and Eq.(17). The solution is obtained by processing $t$ = 2 to $T$ in turn.
Update $\mathbf{P}_{t}$.
This subproblem can be written as
@@
Similar to solving for $\mathbf{A}_{t}$, we can perform SVD on $\mathbf{C}_{t}=\mathbf{Z}_{t}^{T}\mathbf{Z}_{t-1}$ to obtain $\mathbf{P}_{t}=\mathbf{U}_{p}\mathbf{V}_{p}^{T}$.
Update $\mathcal{M}$.
When $\mathbf{A}_{t},\mathbf{Z}_{t},\mathbf{P}_{t},\mathbf{E}_{t}$ are fixed, we can update $\mathcal{M}$ as follow:
$\min_{\mathcal{M}}\beta\|\mathcal{M}\|_{ARMR}+\frac{\rho}{2}\|\mathcal{M}-(\mathcal{Z}+\frac{\mathcal{J}}{\rho})\|_{F}^{2}$ (20)
This is a classic Convex-ConCave Procedure Problem (CCCP) (Yuille and Rangarajan 2001), and we can solve it by the following Lemma:
@@
Let tensor $\mathcal{M}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ with t-SVD. The non-concave tensor rank minimization regularization meets:
$\mathop{\arg\min}\limits_{\mathcal{M}}\frac{\beta}{\rho}\|\mathcal{M}\|_{ARMR}+\frac{1}{2}\|\mathcal{X}-\mathcal{M}\|_{F}^{2}$ (21)
@@
$\tilde{\mathcal{X}}=\Gamma_{\frac{1}{\rho}}(\mathcal{M})=\mathcal{U}*ifft(f(\bar{\mathcal{S}_{f}^{k}})(i,i))*\mathcal{V}^{T}.$ (22)
where ifft denotes the Inverse Fast Fourier Transform. The function $f(\bar{\mathcal{S}_{f}^{k}})$ is a f-diagonal tensor, which is
$f(\bar{\mathcal{S}_{f}^{k}}(i,i),x)=\mathop{\arg\min}\limits_{x\geq 0}\frac{1}{2}(x-\mathcal{S}_{f}^{k}(i,i))^{2}+\left(\frac{1-e^{-x}}{1+e^{-x}}\right)$ (23)
Eq. (23) is a CCCP problem, so we can use the difference of convex programming (Tao and An 1997) to obtain a closed-form solution as:
$f(\mathcal{S}_{f}^{k}(i,i))^{iter+1}=(\mathcal{S}_{f}^{k}(i,i)-\frac{\nabla f(\mathcal{S}_{f}^{k}(i,i))^{iter}}{\rho})_{+}$ (24)
@@
Update $\mathbf{Z}_{t}$.
Fixing the other variables, the optimization function for $\mathbf{Z}_{t}$ can be written as
@@
Updating other variables.
The Lagrange multipliers and penalty parameters can be updated by
@@
$\mu=min(\delta\mu,\mu_{max}),\rho=min(\delta\rho,\rho_{max})$ (28)
where $\rho_{max}$ are the positive constants. The solution of our method is summarized in Algorithm 1, and the convergence analysis is shown in the supplementary material.
Algorithm 1 Updates to subsequent views ($t>1$).
@@
Input: Incremental data $\{\mathbf{X}_{t}\}_{t=1}^{T}$, the latent dimension of basic matrix $m$, cluster number $k$.
Parameters: $\alpha,\beta$, $\lambda$
@@
Output: After obtaining $\mathbf{Z}_{T}$ of the last view, we use $k$-means to obtain the clustering results.
1:Initialize: $\mathbf{A}_{t}$, $\mathbf{P}_{t}$, $\mathbf{E}_{t}$, $\mathbf{Z}_{t}$, $\mathbf{Y}_{t}$ = 0. $\mathcal{Z}$, $\mathcal{J}$, $\mathcal{M}$ = 0, $\mu$ = $\rho=10^{-4}$, $\mu_{max}$ = $\rho_{max}=10^{10}$, $\delta$ = 2.
2:while not converged do
3:  Update $\mathbf{Z}_{hist}$ and $\mathbf{A}_{t}$ by Eq.(6) and Eq. (13).
4:  Update $\mathbf{E}_{t}$ and $\mathbf{P}_{t}$ by Eq. (17) and Eq. (19).
5:  Update $\mathbf{Z}_{t}$ and $\mathcal{M}$ by Eq. (26) and Eq. (24).
6:  Update $\mathbf{Y}_{t}$ and $\mathcal{J}$ by Eq. (27).
7:  Update $\mu$ and $\rho$ by Eq. (28).
8:end while
Complexity Analysis
@@
The computational complexity of our model is analyzed in two stages: the initial view processing ($t=1$) and the subsequent incremental updates ($t>1$). For the initial view, the primary computational costs arise from iteratively updating $\mathbf{A}_{1}$, $\mathbf{Z}_{1}$, and $\mathbf{E}_{1}$. The complexities for these updates are $\mathcal{O}(3dmn+dm^{2})$$\mathcal{O}(dmn)$$\mathcal{O}(dmn)$, respectively. Therefore, the complexity for the initial stage is dominated by $\mathcal{O}(dmn+dm^{2})$ per iteration. For each subsequent incremental step, the complexities for updating $\mathbf{Z}_{hist}$, $\mathbf{P}_{t}$, $\mathcal{M}$, $\mathbf{A}_{t}$, $\mathbf{E}_{t}$, and $\mathbf{Z}_{t}$ are $\mathcal{O}(tdn)$, $\mathcal{O}(m^{2}n+m^{3})$, $\mathcal{O}(m^{3}n+2mntlogt)$, $\mathcal{O}(dmn+dm^{2})$, $\mathcal{O}(dmn)$ and $\mathcal{O}(dmn+m^{2}n)$, respectively. Consequently, the overall complexity per iteration for an incremental step is $\mathcal{O}(tdn+3dmn+m^{3}n+dm^{2}+2m^{2}n+m^{3}+2mntlogt)$. Since $d\ll n$ and $m\ll n$, the computational cost grows approximately linearly with $n$. This linear scalability demonstrates the suitability of our model for processing large-scale datasets.
Experiments
@@
Datasets and Evaluation. To assess the efficacy of the proposed method, we conducted experiments on six benchmark multi-view datasets: ProteinFold, Flower17, GRAZ02, handwritten, Caltech101-all, and MNIST. These datasets exhibit considerable variation in scale, with sample sizes ranging from 694 to 50,000 instances, and the maximum number of views is 27. Besides, three clustering metrics were used for performance evaluation: clustering accuracy (ACC), normalized mutual information (NMI), and adjusted rand index (ARI). Higher value indicates better performance.
Method ProteinFold Flower17 GRAZ02 handwritten Caltech101-all MNIST
@@
FMVACC 27.38 36.50 10.64 49.64 50.27 31.89 52.95 17.55 16.70 91.36 87.65 84.82 23.03 38.93 15.81 53.83 48.98 38.40
AWMVC 34.29 42.93 16.05 56.25 55.34 38.84 37.33 5.91 4.95 74.95 71.77 62.91 27.35 49.65 23.27 54.62 47.03 38.15
3AMVC 29.26 37.33 11.78 45.12 43.30 24.58 39.17 6.01 5.62 90.15 87.05 83.92 22.89 46.05 17.95 38.78 38.58 20.65
NpGC 37.18 43.66 17.32 58.01 57.44 40.34 42.68 8.64 8.17 89.35 80.71 78.10 28.37 49.60 24.17 55.30 47.79 38.77
CAMVC 33.14 42.80 16.67 53.01 50.12 32.77 43.22 13.08 10.22 91.95 84.80 83.11 26.41 47.74 22.18 39.71 36.30 23.34
IWTSN 44.09 70.59 36.27 93.27 92.93 88.79 70.08 34.51 32.91 99.65 99.10 99.22 60.94 87.55 43.85 99.63 98.97 99.18
DSCMC 29.16 39.02 12.80 34.86 44.89 23.59 38.73 5.64 5.17 71.90 68.73 58.76 24.05 42.93 18.00 67.22 59.24 52.35
LMTC 31.63 39.50 14.29 61.38 60.97 44.09 42.35 7.71 7.36 95.10 89.23 89.40 25.20 46.53 39.08 60.34 59.51 48.37
@@
CMVC 31.76 40.55 14.02 55.61 55.02 38.27 48.60 12.22 12.01 89.85 80.25 78.96 24.46 46.88 20.54 60.77 53.30 44.55
LAIMVC 31.89 41.32 15.27 40.37 43.83 25.56 40.88 6.95 6.42 66.98 56.46 48.40 20.34 43.40 16.42 54.37 50.40 39.78
FDMVC 36.51 47.27 19.21 47.17 50.43 31.02 43.78 9.38 8.85 79.68 77.31 69.65 21.98 44.43 18.73 76.79 72.35 66.19
@@
Ours 67.15 82.02 58.15 94.85 94.27 90.56 92.28 80.45 80.52 99.95 99.86 99.89 67.10 89.16 56.97 99.87 99.60 99.73
Table 1: Performance results for six datasets compared to twelve state-of-the-art methods (”OM” stands for Out of Memory).
@@
Comparative Methods. To demonstrate the effectiveness of our methods, we select the following methods for comparison. Traditional Methods: FMVACC (NeurIPS’22) (Wang et al. 2022), AWMVC (AAAI’23) (Wan et al. 2023), 3AMVC(MM’24)(Ma et al. 2024), NpGC (AAAI’24) (Yu et al. 2024), IWTSN (MM’24) (Sun et al. 2024), CAMVC (AAAI’24) (Zhang et al. 2024), DSCMC (PR’25) (Kong et al. 2025), LMTC (CVPR’25) (Liu et al. 2025). Incremental Methods: CMVC (MM’22) (Wan et al. 2022), LAIMVC (MM’24) (Qu et al. 2024), FDMVC (KBS’24) (Lu et al. 2024), FCMVC (TIP’24) (Wan et al. 2024).
@@
Experiment Setup. All experiments were conducted on a personal computer with an Intel Core i9-13900K CPU and 64 GB of RAM. For all comparison methods, we adopted the parameter settings recommended in their original publications. Considering that $k$-means is more sensitive to initialization, we conduct 10 independent experiments and report the average of the clustering metrics to ensure fairness.
Experimental Results and Analysis

### v2

@@ 3.4 The Overall Objective Function
4 Model optimization
We solve Eq.(10) using the Alternating Direction Method of Multipliers (ADMM)  [3]. It decomposes the original problem into easily solvable subproblems by introducing auxiliary variables and Lagrange multipliers. To facilitate the solution, we introduce auxiliary variables $\mathbf{E}_{t}$ and $\mathcal{M}$ instead of $\mathbf{X}_{t}-\mathbf{Z}_{t}\mathbf{A}_{t}$ and $\mathcal{Z}$, respectively. Then, the Lagrangian function can be written as
@@ 4 Model optimization
$+\frac{\mu}{2}\|\mathbf{X}_{t}-\mathbf{Z}_{t}\mathbf{A}_{t}-\mathbf{E}_{t}+\frac{\mathbf{Y}_{t}}{\mu}\|_{F}^{2}+\frac{\rho}{2}\|\mathcal{Z}-\mathcal{M}+\frac{\mathcal{J}}{\rho}\|_{F}^{2},$
where $\mathbf{Y}$ and $\mathcal{J}$ denote the Lagrange multiplier, $\mu>0,\rho>0$ are the penalty parameters.
4.1 Initial view processing ($t=1$).
For the first view, no history information exists, so the objective simplifies to the base view representation learning:
@@
1) Update $\mathbf{A}_{1}$. Fixing $\mathbf{Z}_{1}$, $\mathbf{E}_{1}$, we can obtain $\mathbf{A}_{1}$ by:
@@
This is a standard Orthogonal Procrustes Problem [36]. Let $\mathbf{Q}_{1}=\mathbf{X}_{1}-\mathbf{E}_{1}+\frac{\mathbf{Y}_{1}}{\mu}$. By performing Singular Value Decomposition (SVD) on $\mathbf{Z}_{1}^{T}\mathbf{Q}_{1}$ such that $\mathbf{U}_{A}\mathbf{S}_{A}\mathbf{V}_{A}^{T}=\mathbf{Z}_{1}^{T}\mathbf{Q}_{1}$, the closed-form solution is given by $\mathbf{A}_{1}=\mathbf{U}_{A}\mathbf{V}_{A}^{T}$ (where $\mathbf{U}_{A}$ and $\mathbf{V}_{A}$ are the left and right singular matrices).
2) Update $\mathbf{Z}_{1}$. Fixing the other variables, the subproblem of $\mathbf{Z}_{1}$ can be solved in the following way
@@
3) Update $\mathbf{E}_{1}$. Fixing other variables, $\mathbf{E}_{1}$ can be obtained by
@@
The solution of Eq.(LABEL:E) can be obtained by minimizing the thresholding operator in [42], which is
${(\mathbf{E}_{1})}_{(:,i)}=\left\{\begin{matrix}\frac{\|{(\mathbf{C}_{1})}_{(:,i)}\|_{2}-\frac{1}{\mu}}{\|{(\mathbf{C}_{1})}_{(:,i)}\|_{2}}{(\mathbf{C}_{1})}_{(:,i)},&if\ \|{(\mathbf{C}_{1})}_{(:,i)}\|_{2}>\frac{1}{\mu};\\ 0,&Otherwise.\end{matrix}\right$. (17)
where $\mathbf{C}_{1}=\mathbf{X}_{1}-\mathbf{Z}_{1}\mathbf{A}_{1}+\frac{\mathbf{Y}_{1}}{\mu}$.
4) Update $\mathbf{Y}_{1}$ and $\mu$. The Lagrange multipliers and penalty parameters can be updated by the following rules
$\mathbf{Y}_{1}=\mathbf{Y}_{1}+\mu(\mathbf{X}_{1}-\mathbf{Z}_{1}\mathbf{A}_{1}-\mathbf{E}_{1}),\mu=min(\delta\mu,\mu_{max}),$ (18)
where $\delta>1$ to accelerate convergence, $\mu_{max}$ is the positive constant. $\mathbf{Z}_{1}$ will serve as the starting point and first historical memory for the subsequent optimization process.
4.2 Updates to subsequent views ($t>1$).
1) Update $\mathbf{A}_{t}$ and $\mathbf{E}_{t}$. Similarly, the update rules for $\mathbf{A}_{t}$ and $\mathbf{E}_{t}$ are similar to the first view from Eq.(13) and Eq.(17). The solution is obtained by $t$ = 2 to $V$ in turn.
2) Update $\mathbf{P}_{t}$. This subproblem can be written as
@@
Similar to solving for $\mathbf{A}_{t}$, we can perform SVD on $\mathbf{C}_{t}=\mathbf{Z}_{t-1}\mathbf{Z}_{t}^{T}$ to obtain $\mathbf{P}_{t}=\mathbf{U}_{p}\mathbf{V}_{p}^{T}$.
3) Update $\mathcal{M}$. When $\mathbf{A}_{t},\mathbf{Z}_{t},\mathbf{P}_{t},\mathbf{E}_{t}$ are fixed, we can update $\mathcal{M}$ as follow:
$\min_{\mathcal{M}}\beta\|\mathcal{M}\|_{ARMR}+\frac{\rho}{2}\|\mathcal{M}-(\mathcal{Z}+\frac{\mathcal{J}}{\rho})\|_{F}^{2}.$ (20)
This is a classic Convex-ConCave Procedure (CCCP) [41], and we can solve it by the following Lemma:
@@
Let tensor $\mathcal{M}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ with t-SVD. The non-concave tensor rank minimization regularization is solved by:
$\mathop{\arg\min}\limits_{\mathcal{M}}\frac{\beta}{\rho}\|\mathcal{M}\|_{ARMR}+\frac{1}{2}\|\mathcal{X}-\mathcal{M}\|_{F}^{2}.$ (21)
@@
$\mathcal{X}=\Gamma_{\frac{1}{\rho}}(\mathcal{M})=\mathcal{U}*ifft(f(\mathcal{S}_{f}^{k})(i,i))*\mathcal{V}^{T},$ (22)
where ifft denotes the Inverse Fast Fourier Transform. The function $f(\mathcal{S}_{f}^{k})$ is a f-diagonal tensor, which is
$f(\mathcal{S}_{f}^{k}(i,i),x)=\mathop{\arg\min}\limits_{x\geq 0}\frac{1}{2}(x-\mathcal{S}_{f}^{k}(i,i))^{2}+\left(\frac{1-e^{-x}}{1+e^{-x}}\right).$ (23)
Eq. (23) is a CCCP problem, so we can use the difference of convex programming [31] to obtain closed-form solution:
$f(\mathcal{S}_{f}^{k}(i,i))^{iter+1}=(\mathcal{S}_{f}^{k}(i,i)-\frac{\nabla f(\mathcal{S}_{f}^{k}(i,i))^{iter}}{\rho})_{+}.$ (24)
@@
4) Update $\mathbf{Z}_{t}$. Fixing the other variables, the optimization function for $\mathbf{Z}_{t}$ can be written as
@@
5) Updating other variables. The Lagrange multipliers and penalty parameters can be updated by
@@
$\mu=min(\delta\mu,\mu_{max}),\rho=min(\delta\rho,\rho_{max}),$ (28)
where $\rho_{max}$ is the positive constant. The solution of our method is summarized in Algorithm 1.
Algorithm 1 Updates to subsequent views ($t=2\to V$)
@@
Input: Incremental data $\{\mathbf{X}_{t}\}_{t=2}^{V}$, initial representation $\mathbf{Z}_{1}$, latent dimension $m$, cluster number $k$.
Parameters: $\alpha,\beta$, $\lambda$.
@@
Output: After obtaining last view $\mathbf{Z}_{V}$, we use $k$-means to clustering.
1:for $t=2$ to $V$ do
2:  Initialize: $\mathbf{A}_{t}$, $\mathbf{P}_{t}$, $\mathbf{E}_{t}$, $\mathbf{Z}_{t}$, $\mathbf{Y}_{t}$ = 0. $\mathcal{Z}$, $\mathcal{J}$, $\mathcal{M}$ = 0, $\mu$ = $\rho=10^{-4}$, $\mu_{max}$ = $\rho_{max}=10^{10}$, $\delta$ = 2.
3:  while not converged do
4:   Update $\mathbf{Z}_{hist}$ and $\mathbf{A}_{t}$ by Eq. (6) and Eq. (13).
5:   Update $\mathbf{E}_{t}$ and $\mathbf{P}_{t}$ by Eq. (17) and Eq. (19).
6:   Update $\mathcal{M}$ and $\mathbf{Z}_{t}$ by Eq. (24) and Eq. (26).
7:   Update $\mathbf{Y}_{t}$ and $\mathcal{J}$ by Eq. (27).
8:   Update $\mu$ and $\rho$ by Eq. (28).
9:  end while
10:end for
4.3 Complexity Analysis
@@
Our model comprises two stages: the initial view processing ($t=1$) and the subsequent incremental updates ($t>1$). For the initial view, the primary computational costs arise from iteratively updating $\mathbf{A}_{1}$, $\mathbf{Z}_{1}$, and $\mathbf{E}_{1}$. The complexities for these updates are $\mathcal{O}(3dmn+dm^{2})$, $\mathcal{O}(dmn)$, $\mathcal{O}(dmn)$, respectively. Therefore, the complexity for the initial stage is dominated by $\mathcal{O}(dmn+dm^{2})$ per iteration. For each subsequent incremental step, the complexities for updating $\mathbf{Z}_{hist}$, $\mathbf{P}_{t}$, $\mathcal{M}$, $\mathbf{A}_{t}$, $\mathbf{E}_{t}$, and $\mathbf{Z}_{t}$ are $\mathcal{O}(tmn)$, $\mathcal{O}(m^{2}n+m^{3})$, $\mathcal{O}(m^{2}n)$, $\mathcal{O}(dmn+dm^{2})$, $\mathcal{O}(dmn)$ and $\mathcal{O}(dmn+m^{2}n)$, respectively. Consequently, the overall complexity for an incremental step is $\mathcal{O}(tmn+2dmn+3m^{2}n+dm^{2}+m^{3})$. Since $d\ll n$ and $m\ll n$, the computational cost grows approximately linearly with $n$. This linear scalability demonstrates the suitability of our model for processing large-scale multi-view datasets.
5 Experiments
@@
Datasets and Evaluation. To assess the efficacy of the proposed method, we conducted experiments on six multi-view datasets, details of which are provided in Table 1. Three clustering metrics were utilized for performance evaluation: clustering accuracy (ACC), normalized mutual information (NMI), and adjusted rand index (ARI). Higher values indicate better performance.
Table 1: The details of Incremental Multi-view datasets.
Datasets Samples Clusters Views View Dimensions
ProteinFold 694 12 27 27, …,27
Flower17 1360 17 7 1360, …,1360
GRAZ02 1476 4 6 512,32,256,500,500,680
Handwritten 2000 10 6 216,76,64,6,240,47
Caltech101-all 9144 102 5 48,40,254,512,928
FMNIST 60,000 10 3 1280,512,512
Table 2: Performance results for six datasets compared to thirteen state-of-the-art methods (”OM” stands for Out of Memory).
Methods ProteinFold Flower17 GRAZ02 Handwritten Caltech101-all FMNIST
@@
3AMVC 29.26 37.33 11.78 45.12 43.30 24.58 39.17 6.01 5.62 90.15 87.05 83.92 22.89 46.05 17.95 62.88 49.77 43.00
NpGC 37.18 43.66 17.32 58.01 57.44 40.34 42.68 8.64 8.17 89.35 80.71 78.10 28.37 49.60 24.17 63.37 46.95 41.07
CAMVC 33.14 42.80 16.67 53.01 50.12 32.77 43.22 13.08 10.22 91.95 84.80 83.11 26.41 47.74 22.18 82.57 72.75 66.25
IWTSN 44.09 70.59 36.27 93.27 92.93 88.79 73.71 56.68 52.96 99.65 99.10 99.22 60.94 87.55 43.85 51.58 29.45 26.12
S2MVTC 48.27 72.45 44.48 66.03 82.53 60.41 88.96 70.60 73.14 99.90 99.76 99.78 46.50 76.22 50.89 27.74 11.88 6.57
DSCMC 29.16 39.02 12.80 34.86 44.89 23.59 38.73 5.64 5.17 71.90 68.73 58.76 24.05 42.93 18.00 61.06 40.65 37.57
LMTC 31.63 39.50 14.29 61.38 60.97 44.09 42.35 7.71 7.36 95.10 89.23 89.40 25.20 46.53 39.08 81.70 74.72 70.00
@@
CMVC 31.76 40.55 14.02 55.61 55.02 38.27 48.60 12.22 12.01 89.85 80.25 78.96 24.46 46.88 20.54 OM OM OM
LAIMVC 31.89 41.32 15.27 40.37 43.83 25.56 40.88 6.95 6.42 66.98 56.46 48.40 20.34 43.40 16.42 80.30 76.81 69.75
FDMVC 36.51 47.27 19.21 47.17 50.43 31.02 43.78 9.38 8.85 79.68 77.31 69.65 21.98 44.43 18.73 81.24 78.22 71.19
@@
ASIA 23.92 33.68 8.38 36.76 37.76 20.36 37.13 4.48 3.80 54.85 46.84 36.82 14.30 31.21 7.88 OM OM OM
CDIMVC 37.90 44.93 17.99 40.81 45.26 23.00 41.67 9.69 8.38 79.05 72.74 53.35 20.17 44.37 13.85 OM OM OM
Ours 64.06 80.70 54.97 93.37 94.16 89.56 92.28 80.45 80.52 99.79 99.46 99.53 64.46 87.96 52.27 82.86 76.37 70.80
@@
Comparative Methods. To show the effectiveness of EMIMC, the following methods are selected: 3AMVC [24](ACM MM’24), NpGC [40](AAAI’24), IWTSN [30](ACM MM’24), S2MVTC[20](CVPR’24), CAMVC [43](AAAI’24), DSCMC [18](PR’25), LMTC [19](CVPR’25), CMVC [32](ACM MM’22), LAIMVC [28](ACM MM’24), FDMVC [23](KBS’24), FCMVC [34](TIP’24), ASIA [29](CSVT’25), CDIMVC [10](TKDE’25).
@@
Experiment Setup. All experiments were conducted on a personal computer with an Intel Core i9-13900K CPU and 64 GB of RAM. For all comparison methods, we adopted the parameter settings recommended in their original paper. For our proposed method, we report the average results over five independent runs with different random seeds.
5.1 Experimental Results and Analysis